Hellp me with this.The length a rectangle is 6 inches and its area is (18 x+24) square inches write an expression for the width
Answer by ewatrrr(24785)

 
Hi,
Question States***
Area is 18x+ 24 and L = 6
.
A = Lw The Area Equation gives us the Tool for Solving This.
Area = (18x+24) Question tells us.
Area = 6(3x + 4)= Lw (factoring out the 6 gives us the 'Product" Lw
L = 6 Question tells us
w = (3x+4)